geerky42 (geerky42):

Solve for x: 2sinx + 2cosx = 1

hartnn (hartnn):

for \(a\sin x+b\cos x=c\), divide both sides by \(\sqrt{a^2+b^2}\)

hartnn (hartnn):

then use the formula, for sin Acos B+cos A sin B = sin (A+B)

hartnn (hartnn):

did u try this ?

geerky42 (geerky42):

how can I apply the formula sin Acos B+cos A sin B = sin (A+B)?

hartnn (hartnn):

then \(\cos \alpha =a/\sqrt{a^2+b^2}\) and \(\sin\alpha =b/\sqrt{a^2+b^2}\) so u get \(\cos \alpha \sin x+\sin \alpha \cos x= c/\sqrt{a^2+b^2} \\\sin(x+\alpha )=c/\sqrt{a^2+b^2}\)

OpenStudy (raden):

in right side, that is 1 or 2 ?

geerky42 (geerky42):

1

geerky42 (geerky42):

I'm sorry, I'm lost right here...

hartnn (hartnn):

\(x+\alpha =\sin^{-1}(c/\sqrt{a^2+b^2})\\ x=-\alpha +\sin^{-1}(c/\sqrt{a^2+b^2})\)

hartnn (hartnn):

i've solved for 'x' for any general expression, \(a\sin x+b\cos x=c\) tell me which step you have doubt with ?

geerky42 (geerky42):

But what does α equal to? \(\cos^{-1}{\left(\dfrac{a}{\sqrt{a^2 + b^2}}\right)}\) ?

hartnn (hartnn):

yes, thats correct.

geerky42 (geerky42):

Wow, it looks ugly, but thanks.

hartnn (hartnn):

or arctan(b/a)

geerky42 (geerky42):

ah, much better.

OpenStudy (raden):

if my way : 2sinx + 2cosx = 1 divide both sides by 2 sinx + cosx = 1/2 square both sides sin^2 x + cos^2 x + 2sinxcosx = 1/4 1 + sin2x = 1/4 sin2x = 1/4 - 1 sin2x = -3/4 actually, to get x in right side give a familiar value in sine of special angles that's why i ask before, in right side 1 or 2

geerky42 (geerky42):

but how did you get arctan(b/a) ?

geerky42 (geerky42):

never mind, i got it.

hartnn (hartnn):

if we solve using numbers, it isn't that ugly.... sin alpha/ tan alpha =.... ?

hartnn (hartnn):

squaring both sides may add extraneous roots...